OpenStudy (studentchaitrali):

Let vector u be a constant vector and vector v be a vector of constant magnitude such that magnitude of v is equal to half of magnitude of u and magnitude of u is not equal to zero. Then the max possible angle between vector u and vector(u+v) is: (a) 30 (b) 60 (c) 120 (d) 150

OpenStudy (anonymous):

It might help to take notes of the principles as you read the questions... it really helps because our memory is not meant to memorize too many things at once. we know that u is a constant vector v is the vector of constant magnitude v is half of u and u is never zero.

OpenStudy (anonymous):

I have a feeling that this question may be irrelevant to the answres

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

or \(180^o\) minus that if you want to be pedantic

OpenStudy (mathmate):

 Since |v|/|u|=1/2, we have sin(theta)=1/2, theta=30 degrees. The answers should give degrees as units. They should not be radians.